Assume​ you've generated the following information about the stock of​ Bufford's Burger​ Barns: The​ company's latest dividends of $3.96 a share are expected to grow to ​$4.20 next​ year, to $4.45 the year after​ that, and to $4.72 in year 3. After​ that, you think dividends will grow at a constant 5​% rate.

a. Use the variable growth version of the dividend valuation model and a required return of 15​% to find the value of the stock.

b. Suppose you plan to hold the stock for three​ years, selling it immediately after receiving the $4.72 dividend. What is the​ stock's expected selling price at that​ time? As in part a​, assume a required return of 15​%.

c. Imagine that you buy the stock today paying a price equal to the value that you calculated in part a.You hold the stock for three​ years, receiving dividends as described above. Immediately after receiving the third​ dividend, you sell the stock at the price calculated in part b. Use the IRR approach to calculate the expected return on the stock over three years. Could you have guessed what the answer would be before doing the​ calculation?

d. Suppose the​ stock's current market price is actually ​$40.71. Based on your analysis from part a​,is the stock overvalued or​ undervalued?

e. A friend of yours agrees with your projections of​ Bufford's future​ dividends, but he believes that in three​ years, just after the company pays the ​$4.72​dividend, the stock will be selling in the market for ​$55.10. Given that​ belief, along with the​ stock's current market price from part d​, calculate the return that your friend expects to earn on the stock over the next three years.
Answer #1

Answer:

A)

PV of Stock = 3.96+4.20/(1+15%)+4.45/(1+15%)^2+4.72/(1+15%)^3+4.72*(1+5%)/(15%-5%)=$63.64

B)

Stock expected selling price after three years= 4.72*(1.05)/(15%-5%)=$49.56

C)

Cash flow

Year Dividend Selling price Total
0 3.96 3.96
1 4.2 4.2
2 4.45 4.45
3 4.72 49.56 54.28

let IRR is equal to x% so at IRR NPV=0

NPV= PV of all dividends+pv of selling price - initial payment for Stock

initial payment for Stock=PV of all dividends +pv of selling price

63.64=3.96+4.20/(1+x%)+4.45/(1+x%)^2+4.72/(1+x%)^3+49.56/(1+x%)^3

Solving for x=1.92%

D) Since current stock price is $40.71 which is less than value of stock ($63.64) so stock is undervalued.